OpenStudy (juan1857):

Find the point of inflection of y=2sin(x)+cos(2x) ;[0,pi]

OpenStudy (juan1857):

I already found the second derivative which is \[-2\sin(x)-4\cos(2x)\]

OpenStudy (juan1857):

I am just having trouble finding when is it 0

OpenStudy (solomonzelman):

Any \(x=\alpha\) that satisfies \(f''(\alpha)=0\), as long as the signs on the sides of \(f''(\alpha)\) are opposite (so that for a sufficiently small \(\delta\), you have \(f''(\alpha-\delta)\) and \(f''(\alpha+\delta)\) with opposite signs).

OpenStudy (solomonzelman):

I will say this in a better less abstract term.

OpenStudy (mathmale):

Just set the 2nd derivative = to 0 and solve the resulting equation for x. Can you do that?

OpenStudy (solomonzelman):

Suppose that you have a function \(f\), and \(f''(3)=0\), then \(f''(3.04)\) and \(f''(2.96)\) have opposite signs.

OpenStudy (solomonzelman):

(If the signs of \(f''(3.04)\) and \(f''(2.96)\) are not opposite, - i.e. both negative or both positive, then \(x=3\) is not an inflection point.)

OpenStudy (solomonzelman):

So, at first you have to find the \(x\) values, (the set of \(x\) values) that satisfies \(f''(x)=0\).

OpenStudy (solomonzelman):

\(-2\sin(x)-4\cos(2x)=0\) \(\sin(x)=-2(1-2\sin^2(x))\) \(4(\sin x)^2-\sin(x)-2=0\) proceed with a substitution

OpenStudy (solomonzelman):

(or with questions if you have them)

OpenStudy (juan1857):

I am lost on how you got sinx on one side by itself @SolomonZelman

OpenStudy (solomonzelman):

\(-2\sin(x)-4\cos(2x)=0\) \(\color{red}{-2\sin(x)=4\cos(2x)}\) \(\color{red}{\sin(x)=-2\cos(2x)}\) \(\sin(x)=-2(1-2\sin^2(x))\) \(\color{red}{\sin(x)=-2+4\sin^2(x)}\) \(4(\sin x)^2-\sin(x)-2=0\) the steps in red are the once I omitted (due to thinking they are obvious)

OpenStudy (juan1857):

hmmm

OpenStudy (juan1857):

I understand how you divided by 2

OpenStudy (juan1857):

-2

OpenStudy (solomonzelman):

you don't possess a firm knowledge of algebra?

OpenStudy (juan1857):

Its been quite a while

OpenStudy (solomonzelman):

Algebra, as a matter of fact, is applied constantly to any math problem ... (Well, maybe that is just my experience alone)

OpenStudy (solomonzelman):

In any case, can you solve the equation from there?

OpenStudy (juan1857):

so would it also be cos(2x)=sin(x)

OpenStudy (solomonzelman):

sin(x)=−2cos(2x)

OpenStudy (solomonzelman):

I have already done the algebra up to some point ... sub z=sin(x)

OpenStudy (solomonzelman):

and solve the quadratic equation.